OpenStudy (anonymous):

Hunter drove 27 mi in 1/2 h. How far will he drive in 2 1/2 h at the same speed? ? mi

OpenStudy (anonymous):

How many 0.5hrs are there in 2.5hrs? Multiply this number by the distrance he travelled in half an hour to find the answer.

OpenStudy (anonymous):

this is how you will do it add 27 each half our passes like 27*5=135

OpenStudy (anonymous):

do i get a medal???? PLZ

OpenStudy (anonymous):

do I divied it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you add 27 to each half an hour passes to get your awnser :D

OpenStudy (anonymous):

and i told you the awnser so :D:D:D:D:D

OpenStudy (anonymous):

what is the answer though?

OpenStudy (anonymous):

som1 reply

OpenStudy (anonymous):

@LondonRulerOfEngland We know he travels 27 miles in half an hour. All we have to do is scale this up: 0.5 hrs = 27 miles 1 hr = 54 miles 1.5hrs = 81 miles 2 hrs = 108 miles 2.5 hrs = 135 miles You could do this a shorter way: \[\frac{2.5}{0.5}*27=5*27=135\]

OpenStudy (anonymous):

soo my answer is 135